OpenStudy (anonymous):

If a=k(m)^-1 and b=k(m)^1

OpenStudy (anonymous):

here a=k/m and b=km ... do u need ab =k^2 or else ur quest????

OpenStudy (anonymous):

Nope, thanks! :]